OpenStudy (anonymous):

Given the following functions f(x) and g(x), solve (f + g)(3) and select the correct answer below: f(x) = 2x + 21 g(x) = x − 24 A)−27 B)−21 C)6 D)48 I THINK ITS D BUT IM NOT SURE HELP ME PLEASE IM RIGHT?

OpenStudy (anonymous):

@chrisdbest

OpenStudy (anonymous):

@Haseeb96 @horsegirl27

OpenStudy (haseeb96):

just add these two equations as we add two numbers . and show me what you get

OpenStudy (chrisdbest):

No, you have to plug in 3 for ALL X Values

OpenStudy (chrisdbest):

So f(x) = 27 & g(x) = -21

OpenStudy (chrisdbest):

So 27 + -21 = Your answer

OpenStudy (anonymous):

I HAVE -48

OpenStudy (haseeb96):

you will get (f+g)(x) = 2x+21 +x-24 = 3x -3 now put x=3 in order to find (f+g)(3) (f+g)(3) = 3(3) -3 = 9-3 = 6 is your answer

OpenStudy (anonymous):

OH I GET NOW THANK YOU

OpenStudy (anonymous):

HEY, CAN I ASK YOU WERE DID YOU GET 9? @Haseeb96 ??

OpenStudy (haseeb96):

first i put 3 at the place of x and then multiply 3 with 3

OpenStudy (anonymous):

YEA I GET NOW THANKS